### v1

# LOSC: LiDAR Open-voc Segmentation Consolidator

Abstract
We study the use of image-based Vision-Language Models (VLMs) for open-vocabulary segmentation of lidar scans in driving settings. Classically, image semantics can be back-projected onto 3D point clouds. Yet, resulting point labels are noisy and sparse. We consolidate these labels to enforce both spatio-temporal consistency and robustness to image-level augmentations. We then train a 3D network based on these refined labels. This simple method, called LOSC, outperforms the SOTA of zero-shot open-vocabulary semantic and panoptic segmentation on both nuScenes and SemanticKITTI, with significant margins.
Keywords: LiDAR segmentation, LiDAR semantic segmentation, LiDAR panoptic segmentation, 3D scene understanding, VLM, open-vocabulary, zero-shot
002020202040404040nuScenes                      SemanticKITTIOV3D [19]GGSD [16]AFOV [17]LOSC (ours)-3.8-3.2-1.449.3SAL [22]LOSC (ours)35.2-6.5
Semantic segmentation (mIoU%)
002020202040404040nuScenes              SemanticKITTISAL [22]LOSC (ours)-10.048.4SAL [22]SAL-4D [24]LOSC (ours)-7.6-1.632.4
Panoptic segmentation (PQ%)
Figure 1: Performance of LOSC compared to SOTA on zero-shot open-voc segmentation.
1 Introduction
Vision-Language Models (VLMs) are flourishing, offering unprecedented capabilities [1, 2, 3, 4]: employed as general Vision Foundation Models (VFMs), they are usable off-the-shelf for many open-vocabulary tasks: image classification, visual reasoning, visual question answering, image retrieval, image captioning, object detection, semantic segmentation, panoptic segmentation, etc. However, for a number of closed-set and domain-specific tasks, including object detection and semantic or instance segmentation, zero-shot performance of VLMs is not up to the level of supervised finetuning with manual labels [5]. Several reasons may explain it.
First, some VLMs are trained with the sole supervision of image-text pairs, which makes dense tasks such as segmentation harder to learn [6]. Second, as training a VLM requires a huge quantity of data, some models are trained on data scrapped from the Web, which are thus subject to a distribution bias: they contain comparatively few samples from particular domains, e.g., medical images [7, 8]. This concerns both the visual contents (images) and the associated vocabulary (text). Third, “vision” in “VLM” generally only implies images, as 3D data (depth maps, point clouds, etc.) aligned with text is in much shorter supply than 2D data [9].
Therefore, most existing 3D VLMs (relating 3D data and text) actually operate via intermediate 2D data, leveraging a 2D VLM already trained on a large quantity of images: the 2D VLM is used to extract 2D information (labels or features), which is then back-projected onto 3D [9]. The 3D labels (or features) can then be directly employed as such, or used as pseudo-labels to train a 3D model, as done for the feature-based distillation of a 2D VFM into a 3D VFM [10].
In this paper, we study the use of 2D VLMs for the zero-shot open-vocabulary segmentation of lidar scans in driving settings, transferring 2D knowledge to 3D while specializing it for a specific purpose. Precisely, given (i) an open-vocabulary 2D VLM, (ii) any set of classes, defined via arbitrary textual prompts, and (iii) an unannotated 3D dataset of lidar scans with registered images, we create a 3D model that efficiently can infer the semantic segmentation of any lidar scan for these specific classes. Although this 3D model is closed-set, its specification, based on textual prompts, is fully open-set, and the process of creating the 3D model is annotation-free. Besides, as we can leverage a state-of-the-art (SOTA) 3D segmenter, which we train from the segments produced by the 2D VLM, we obtain an efficient 3D model.
However, driving scenes suffers from the data shortage mentioned above: for this specific domain, there is relatively little visual content available associated with text [7, 8], and even less involving 3D data [9]. As a result, current VLMs still lag behind full supervision for the semantic and panoptic segmentation of driving scene images, and even more so for lidar scans [5, 11, 12].
Concretely, in driving scenes, VLM-based image segments are often missing or wrongly labeled, while object boundaries are imprecise. This label inaccuracy transfers to 3D point after back-projection. Additionally, some 3D points may get wrong 2D information due to parallax between camera and lidar sensors, and 3D points invisible from cameras (not in their field of view) do not get any information.
We propose three label consolidation mechanisms to address these issues. First, we consider that semantics should be robust to basic image augmentations and discard labels that change according such variations. Second, we assume that 3D labeling should be consistent over time while driving, at least regarding the static part of the scene. Therefore, we ensure that all points falling into the same voxel get the same label over a whole driving sequence. Third, it is known that iterated training has a denoising effect on pseudo-labels. Therefore, starting from a (self-supervised) pretrained 3D model, we run several finetuning iterations, interleaving them with spatio-temporal consolidation.
To get the best performance, we additionally (i) benchmark existing VLMs for semantic segmentation, (ii) experiment with different prompt variants, and (iii) leverage a pretrained 3D backbone. It allows us to improve the state of the art (SOTA) of open-vocabulary 3D semantic segmentation by a significant margin on two classic datasets (see Fig. 1). Our contributions are as follows:
  • We propose a way to combine several techniques to improve VML-based 3D segmentation, including augmentation- and time-based consolidations, 3D model pretraining and iterated finetuning.
  • We additionally study the specific choice of a basis VLM for the driving setting, for both semantic and panoptic segmentation, using different metrics and prompts.
  • Despite the simplicity of our method (no complex training or architecture, no prompt engineering or VLM tweaking, very few and little-sensitive parameters), we largely improve the SOTA of open-vocabulary semantic and panoptic segmentation of lidar data.
2 Related work
Open-vocabulary 2D segmentation.
Models such as CLIP [6] and ALIGN [13], which aligns global text and image representations, demonstrate strong capabilities in zero-shot classification. These capabilities have been extended to dense prediction tasks requiring pixel-level information, such as semantic segmentation. Some methods leverage the original CLIP model and find recipes to extract dense pixel representations from it. For example, MaskCLIP [14] extract patch features in the last attention layer. CLIP-DIY [15] aggregates features from several crops. Some other methods finetune or train from scratch a CLIP-like architecture but where text representations are aligned with pixel- or patch-level representations [16, 17, 18, 19, 20, 21]. In this work, we evaluate the following models: OpenSeg [17], SEEM [18], OpenSeed [19] and Grounded-SAM [20].
Open-vocabulary 3D segmentation.
The capabilities of open-vocabulary 2D semantic segmentation can be transferred to 3D network by distillation.
OpenScene [22] and CLIP2Scene [23] do so by leveraging point-pixel correspondences and training a 3D network to produces point features which aligns with the corresponding pixel representations extracted from a CLIP-like model. The process was then improved by exploiting, e.g., additional geometric regularization [24, 25], improving text queries [26, 27], or exploiting models such as SAM [20] for denoising [28, 29].
SAL [30] leverages SAM [20], to get instance masks from images, and MaskCLIP [31], to get per-mask CLIP features. These masks and features are lifted in 3D (thanks to point-pixel correspondences) and a 3D network is trained to predict these pairs of masks and features. SAL therefore enables zero-shot lidar panoptic segmentation. SAL was recently improved by operating on sequences rather than single scans in [32]. In this work, we train our 3D network for semantic segmentation and leverage ALPINE [33] to obtain panoptic labels.
LeAP [34], developed concurrently, is closer to our method. LeAP leverages a VLM to generate 2D soft labels, which are lifted in 3D. The 3D soft labels are refined using a Bayesian update leveraging time consistency. To further improve the quality of these 3D soft labels, a 3D network is trained on the most confident soft labels and the output of this trained 3D network is used to enrich the 3D soft labels originating for images. Unlike LeAP, we exploit one-hot labels instead of soft-labels and a simple majority voting when leveraging time-consistency, instead of a Bayesian update. But more importantly, we provide new insights by (a) comparing the performance of several VLMs, (b) showing how to best combine image augmentations and time-consistency to improve pseudo-label quality, and (c) demonstrating that using a pretrained 3D network gives a significant performance boost.
3 Method
Figure 2: Overall pipeline of LOSC. The images are first segmented using an open-vocabulary segmentation model. The semantic 2D labels are then backprojected on the lidar points. These 3D labels are refined using three consolidation steps: the first uses label consistency across several image augmentation, the second uses time consistency, the last combines the best labels from the previous steps. These labels are used to finetune a 3D network. Finally, few steps of self-training are used to improve the results.
Our goal is the zero-shot open-voc labeling of lidar scans. We assume we are given (i) a closed but unconstrained set of classes $\mathcal{C}$, (ii) a VLM providing semantic segments of images from open-vocabulary textual prompts, (iii) an unlabeled training dataset $\mathcal{D}$ of sequences of synchronized and calibrated images and scans. The overall pipeline of our method, illustrated on Fig. 2, is as follows:
  1. 1.
    Based on prompts for the target set of classes, we use the VLM to produce a semantic segmentation of the dataset of images. Back-projecting these 2D segments onto the associated point clouds then provides initial, basic semantic labels for 3D points.
  2. 2.
    These initial 3D semantic labels are first consolidated using a temporal aggregation of lidar scans in a sequence, and voxel-level majority voting.
  3. 3.
    Considering a set of augmentations applied to images before they go through the VLM, we consolidate 3D point labels by only keeping those that remain consistent.
  4. 4.
    Next, we combine augmentation-based consolidation and time-based consolidation, ensuring that label classes are represented enough to be used for training.
  5. 5.
    Finally, the resulting labels are used to finetune a pretrained 3D network. Iterating a few times temporal voxel consolidation and finetuning further improves segmentation quality.
Each of these steps is detailed in the following subsections.
3.1 Initial Zero-shot Semantic Segmentation of LiDAR data with the VLM
Formally, given an image $I$ and textual prompts $T_{\mathcal{C}}$ for the target classes $\mathcal{C}$, the VLM produces a 2D pixelwise semantic segmentation $S_{I}=\text{VLM}(I,T_{\mathcal{C}})$. It is back-projected into the 3D points of the associated lidar point cloud $P$ to provide a 3D pointwise semantic segmentation $S_{P}=\Pi_{I,P}(S_{I})$, where $\Pi_{I,P}$ denotes the back-projection. 3D points falling outside of the camera frustrum, or hidden behind other points due to parallax configurations between camera and lidar, get the ignore label. When a lidar scan $P$ is captured together with several images $\mathcal{I}(P)$, the images are segmented and back-projected jointly, forming the initial VLM-based point cloud labeling $L_{\text{vlm}}$:
$L_{\text{vlm}}^{P}$ $=\cup_{I\in\mathcal{I}(P)}\;\Pi_{I,P}(\text{VLM}(I,T_{\mathcal{C}}))$ (1)
(In the following, we drop subscripts, superscripts and parameters if they are clear from the context.) These initial labels are both noisy (incorrect VLM labeling, camera-lidar parallax inconsistencies) and sparse (missing 2D VLM labeling, 3D points invisible from cameras) in the sense that a point $p\in P$ is given a label $L_{\text{vlm}}(p)\in\mathcal{C}\cup\{\emph{ignore}\}$, where ignore represents the lack of semantics. For a dataset like SemanticKITTI [35], this sparsity is particularly severe (cf. Section 4), because the 360 lidar is associated to only one camera, with a 90 field of view. In the following, we introduce three refinement strategies to improve both label quality and quantity.
3.2 Label Refinement with Time-Based Consolidation (TBC)
Due to resolution limit, faraway objects are hard to identify because they are perceived with only few pixels and points. However, as the ego vehicle drives and comes closer to them, their semantics becomes more apparent. To leverage semantic consistency over time, we aggregate each sequence $\mathcal{P}$ of labeled point clouds into the same coordinate system and enforce label consistency at voxel level by majority voting. Then, for any individual scan $P\in\mathcal{P}$, we assign to each point $p\in P$ the label of the voxel $p$ falls into. It produces scans $L_{\text{tim}}^{\mathcal{P}}=\text{TBC}(L_{\text{vlm}}^{\mathcal{P}})$ with time-consistent labels.
This process may increase label sparsity because, in our implementation, ignore is treated as a full-fledged class label. The idea is that a majority of points in a voxel vote for ignore, then all points falling into that voxel should be considered as unreliable and labeled as ignore too. Besides, even when such a configuration does not arise, the label distribution may also change, with some classes loosing point representatives in favor of other classes.
A variant consists in weighting the label votes with the softmax values of the label prediction. However, empirically, the gain is only marginal and it is simpler just to count class labels in voxels.
3.3 Label Refinement with Augmentation-Based Consolidation (ABC)
A robust network should be little sensitive to small input variations. To filter out “brittle” labels, we probe VLM labeling confidence using image augmentations: a label is considered as untrustful if it changes across a set of image augmentations. This Augmentation-Based Consolidation (ABC) can be seen as a form of Test-Time Augmentation (TTA) applied to VLM inference.
Formally, we consider a set of image augmentations $\mathcal{A}$. For any augmentation $A\in\mathcal{A}$, we denote by $L_{\text{vlm}}^{A}$ the initial point cloud labels produced by the VLM from augmented images $A(\mathcal{I}(P))$. If the label of a point $p\in P$ is the same class $c$ whatever the augmentation $A\in\mathcal{A}$, then the augmentation-based consolidated label $L_{\text{aug}}(p)$ of $p$ is this class $c$; otherwise, it is the ignore label:
$L_{\text{vlm}}^{A}$ $=\cup_{I\in\mathcal{I}}\;\Pi_{I}(\text{VLM}(A(I),T_{\mathcal{C}}))$ (2)
$L_{\text{abc}}(p)$ $=\text{ABC}(L_{\text{vlm}}^{\mathcal{A}})(p)=\begin{cases}c&\text{if~{}~{}}\forall A\in\mathcal{A},L_{\text{vlm}}^{A}(p)=c\\ \emph{ignore}&\text{otherwise}\end{cases}$ (3)
We then enforce temporal voxel consistency and build a consolidated labeling $L_{\text{aug}}=\text{TBC}(L_{\text{abc}})$.
3.4 Label Refinement Combining Augmentation- and Time-based Consolidation (ATC)
Even more than time-based consolidation, augmentation-based consolidation tends to increase label quality but also to reduce label quantity (cf. Section 4). To make sure points of each class remain both numerous enough and balanced enough, we introduce a way to combine $L_{\text{aug}}$ and $L_{\text{tim}}$.
Let $N^{c}=|\{L(p)=c\}_{p\in P}|$ be the number of points in a labeling $L$ of point cloud $P$ that have been given label $c$. The set of classes in $\mathcal{C}$ that are robust to augmentation-based consolidation is
$\mathcal{C}_{\text{robust-aug}}$ =$=\{c\in\mathcal{C}\mid N_{\text{aug}}^{c}\geq N\text{~{}and~{}}N_{\text{aug}}^{c}/N_{\text{tim}}^{c}\geq\tau\}$ (4)
for some minimum number of points $N$ and minimum proportion of preserved points $\tau$. For these classes, it is safe to use the augmentation-based labeling $L_{\text{aug}}$. Otherwise, it is safer to stick to time-based labeling $L_{\text{tim}}$. Therefore, we construct the following combined labeling $L_{\text{aug}{\otimes}\text{tim}}$:
[…]
This hybrid strategy helps preserve class balance and label quality, especially for rare classes and for categories that are hard to predict.
3.5 Label Refinement with Iterative FineTuning (FT)
Finally, we use the combined consolidated labels of $L_{\text{aug}{\otimes}\text{tim}}$, obtained on the whole training set $\mathcal{D}$, as pseudo-labels to train a 3D network. While all previous labelings ($L_{\text{vlm}}$, $L_{\text{tim}}$, $L_{\text{aug}}$, $L_{\text{aug}{\otimes}\text{tim}}$) are partial, in the sense that some points may be labeled as ignore, applying the trained 3D network on any scan yields a fully-labeled point cloud (no ignore).
To improve the performance of this 3D segmenter, we do two things. First, we actually start from a pretrained model $M_{0}$ and finetune it (FT) on the pseudo-labels of $L_{\text{aug}{\otimes}\text{tim}}(\mathcal{D})$, resulting in a model $M_{1}=\text{FT}_{L_{\text{aug}{\otimes}\text{tim}}(\mathcal{D})}(M_{0})$. Second, we iterate training, as often done in semi-supervised settings. More precisely, given a trained model $M_{n}$, we create new pseudo-labels $M_{n}(\mathcal{D})$ for the training set, apply time-based consolidation, and use these refined pseudo-labels to finetune $M_{n}$, resulting a new model $M_{n+1}=\text{FT}_{\text{TBC}(M_{n}(\mathcal{D}))}(M_{n})$. Please note that augmentation-based consolidation (ABC) is performed only once and used via $L_{\text{aug}{\otimes}\text{tim}}$ to finetune $M_{1}$. For all subsequent iterations, only time-based consolidation is used, to refine pseudo-labels of the training set before the next finetuning. The resulting model, after a few iterations, is our final model, which we refer to as LOSC.
4 Experiments
Method Segmentation Prompt Label coveragemAccmIoU%
Semantic Panoptic Minimal Rich
Grounded SAM [36] 68.0 32.8 21.1
Grounded SAM 2 [36] 75.3 46.6 24.3
OpenSeeD [19] 74.9 43.4 33.4
74.8 42.8 33.8
63.9 39.5 31.9
66.9 39.9 32.4
SEEM [18] 76.2 47.7 29.6
76.3 44.7 25.7
39.9 25.2 21.1
23.3 13.1 8.3
OpenSeg [17] 72.8 46.5 26.7
76.3 43.7 24.0
Table 1: Semantic segmentation benchmark of 2D open-vocabulary VLMs on the nuScenes validation set. Performance is evaluated using simple or rich prompts, with semantic or (merged) panoptic segmentation results where applicable.
In this section, we analyze our proposed method and compare it against the current SOTA.
Implementation details.
On the image side, for augmentation-based consolidation, we use the Albumentations library [37], applying 10 distinct image transformations: horizontal-flip, hue-saturation, blur, color-jitter, auto-contrast, sharpen, chromatic-aberration, emboss, fancy-pca, clahe. The appendix details the rationale of this choice. This set of augmentations is consistent in the sense that although a single disagreement is enough to discard the label of a point (cf. Eq. 3), empirical results show the robustness of pseudo-labeling (cf. appendix), leading to SOTA results (cf. Table 5).
Regarding 3D model pretraining, we leverage a self-supervised backbone to benefit from class-agnostic representations for our open-set segmentation approach. For our 3D network, we thus chose WaffleIron [38] (as did LeAP [34]), more precisely WI-48-768, for which self-supervised pretrained weights with ScaLR [10] are publicly available. This pretrained model is currently among SOTA self-supervised 3D backbones for lidars. When finetuning this network, we use the protocol of [10] but reduce the number of epochs to 10. For trainings without ScaLR initialization, we train the model from scratch as described in [38].
For spatio-temporal consolidation, we set the voxel size to 10 cm. For the combination of consolidations, we set $N=200k$ and $\tau=1/3$. Last, we limit ourselves to 3 finetuning iterations, which defines our method, LOSC. We experiment on both nuScenes [39] and SemanticKITTI [35] with identical parameters.
Benchmarking 2D VLMs.
We evaluate the quality of 3D semantic labels $L_{\text{vlm}}$ obtained from several established 2D open-vocabulary VLMs after back-projection: Grounded SAM [36], SEEM [18], OpenSeeD [19] and OpenSeg [17]. We use two types of prompts in our experiments: minimal and rich prompts (which are detailed in the Appendix). Some VLMs can output either semantic or panoptic masks. We tested both alternatives: panoptic masks were turned into semantic masks by simply merging all instance-masks from each class. The quality of $L_{\text{vlm}}$ labels is evaluated on the validation set on nuScenes. We compute the label coverage (percentage of annotated points after back-projection of annotated pixels) as well as the mAcc and mIoU, where non-annotated points after back-projection are counted as errors. The results are presented in Table 1.
First, we notice that about 65 to 75% of the points receive a label in most configurations, with the exception of the panoptic variant of SEEM where less than 40% of the points get a label. The differences in label coverage is explained by the fact that the VLMs do not provide labels for some pixels. Second, our results show moderate performance differences between the minimal and rich prompts, especially for OpenSeeD, but except for SEEM. For the Grounded SAM family of models, we report results only with minimal prompts as the runtime increases substantially with more elaborate prompts. Third, we notice that the semantic variant of VLMs tends to provide labels of better quality than the (merged) panoptic variant. Finally, we notice that OpenSeeD in the best performing model regarding mIoU. Based on these results, we generate $L_{\text{vlm}}$ for the training sets of SemanticKITTI and nuScenes by back-projecting 2D labels obtained using the semantic variant of OpenSeeD with minimal prompts, i.e., the fastest, simplest and nearly best discovered configuration.
3D model pretraining and label refinement.
Time Augmentation ScaLR nuScenes SemanticKITTI
Labeling consistency consistency pretraining mIoUmIoU%
$L_{\text{vlm}}=\Pi(\text{VLM}(\mathcal{I},T_{\mathcal{C}}))$ 37.4 06.5
40.5 22.6
$L_{\text{tim}}=\text{TBC}(L_{\text{vlm}})$ 46.4 32.1
$L_{\text{abc}}=\text{ABC}(L_{\text{vlm}})$ 39.2 23.0
$L_{\text{aug}}=\text{TBC}(\text{ABC}(L_{\text{vlm}}))$ 46.5 31.1
$L_{\text{aug}{\otimes}\text{tim}}=\text{comb}(L_{\text{aug}},L_{\text{tim}})$ 45.7 31.0
48.0 34.2
Table 2: Effect of our label consolidations on nuScenes and SemanticKITTI validation sets.
Table 2 shows the benefits of leveraging a pretrained 3D network and of using our label refinement strategies.
First, we observe that training from ScaLR pretrained weights, compared to random weights, significantly boosts performance by 3.1 and 16.1 mIoU points on nuScenes and SemanticKITTI, respectively, when starting from $L_{\text{vlm}}$. Starting from pretrained weights continues to improve performance even when starting from our most refined pseudo-labels $L_{\text{aug}{\otimes}\text{tim}}$.
Second, applying our time-based consolidation strategy (TBC) directly on $L_{\text{vlm}}$, producing $L_{\text{tim}}$, further improves performance by about 6 and 10 points on nuScenes and SemanticKITTI, respectively.
Third, our augmentation-based consolidation (ABC), directly applied on $L_{\text{vlm}}$ to produce $L_{\text{abc}}$, brings little gain on SemanticKITTI over just training from $L_{\text{vlm}}$, and even slightly degrades performance on nuScenes. Besides, refining $L_{\text{abc}}$
with TBC to produce $L_{\text{aug}}$ leads to similar results than with just $L_{\text{tim}}$. But ABC is actually not meant to operate on it own. The benefit of augmentation-based consolidation appears in fact when combining $L_{\text{aug}}$ and $L_{\text{tim}}$, where we obtain an extra boost of performance of 1.6 and 2.1 mIoU points on nuScenes and SemanticKITTI, respectively, compared to using $L_{\text{tim}}$ alone.
Iterative finetuning.
The results in Table 3 show that the quality of labels keeps improving after each round of finetuning, whether starting from $L_{\text{tim}}$ labels or $L_{\text{aug}{\otimes}\text{tim}}$. The best results are obtained when starting from $L_{\text{aug}{\otimes}\text{tim}}$, which shows that our iterative finetuning strategy is able to maintain the initial advantage provided by our most-refined labels. While a single iteration may provide a gain up to +2.4 mIoU pts (first iteration on SemanticKITTI), further iterating provides diminishing returns. What thus makes sense practically is to only iterate a few times. In LOSC, we perform 3 finetuning iterations.
nuScenes  (mIoU%) SemanticKITTI  (mIoU%)
Labeling 1st iter. 2nd iter. 3rd iter. 1st iter. 2nd iter. 3rd iter.
$L_{\text{tim}}=\text{TBC}(L_{\text{vlm}})$ 46.4 47.3 47.7 32.1 33.0 33.2
$L_{\text{aug}{\otimes}\text{tim}}=\text{comb}(L_{\text{aug}},L_{\text{tim}})$ 48.0 48.9 49.3 34.2 35.0 35.2
Table 3: Effect of iterative finetuning on nuScenes and SemanticKITTI validation sets.
nuScenes SemanticKITTI
Labeling Iter. Label coverage % mIoU % Label coverage % mIoU %
$L_{\text{vlm}}=\Pi(\text{VLM}(\mathcal{I},T_{\mathcal{C}}))$ 0 75.8 39.3 15.4 25.5
$L_{\text{tim}}=\text{TBC}(L_{\text{vlm}})$ 0 74.1 39.9 6.7 26.3
$L_{\text{abc}}=\text{ABC}(L_{\text{vlm}})$ 0 63.7 45.1 11.9 28.1
$L_{\text{aug}}=\text{TBC}(\text{ABC}(L_{\text{vlm}}))$ 0 61.6 46.1 04.9 29.4
$L_{\text{aug}{\otimes}\text{tim}}=\text{comb}(L_{\text{aug}},L_{\text{tim}})$ 0 61.9 46.6 05.0 31.4
$L_{\text{aug}{\otimes}\text{tim}}=\text{comb}(L_{\text{aug}},L_{\text{tim}})$ 1 100.0 46.8 100.0 33.6
$L_{\text{aug}{\otimes}\text{tim}}=\text{comb}(L_{\text{aug}},L_{\text{tim}})$ 2 100.0 47.5 100.0 34.4
Table 4: Quantity and quality of training labels computed on nuScenes and SemanticKITTI training sets.
Pseudo-label analysis.
In this section, we analyze both the quantity and quality the data annotations used for training. The results are presented in Table 4 for nuScenes and SemanticKITTI. We observe that the VLM struggles more (lower mIoU) on SemanticKITTI than on nuScenes. Additionally, with each level of label refinement, the number of labeled points decreases, but the quality of the labels improves. Thanks to our iterative finetuning framework, we are able to annotate the entire datasets after the first iteration while maintaining a high annotation quality.
Comparison to state-of-the-art zero-shot semantic segmentation.
[…]
LOSC outperforms the SOTA by +3.2 mIoU pts on nuScenes. It even outperforms methods that additionally use images at inference time by +1.4 mIoU point. The gain over the SOTA on SemanticKITTI is +6.5 mIoU points (but only a couple of other methods evaluate on this dataset). Note that we could not compare against LeAP [34], which evaluates with a specific protocol that is not fully specified. More specifically, the class mapping from the original and classical 19 classes of SemanticKITTI to their 11 classes is not provided, and the code is not publicly available.
Compared to some other approaches, we obtain those SOTA results without resorting to any prompt engineering, VLM tweaking, intricate architecture, complex training, or the use of images for inference. Besides, we only have very few and robust parameters, which are the same for all datasets. For instance, we set $\tau$ to 1/3 but any value between 1/4 and 1/2 would have given equal or similar results for both nuScenes and SemanticKITTI. Beyond these bounds, performance would start dropping a bit.
Method nuScenes SemanticKITTI
WaffleIron (full supervis.) [38] 78.7 63.4
CLIP2Scene [23] 20.8 -
Towards VFMs [40] 26.8 -
AutoVoc3D w. LAVE [26] 30.6 -
AdaCo [29] 31.2 25.7
SAL [30] 33.9 28.7
OV3D w/ OpenScene-3D [27] 45.5 -
GGSD [24] 46.1 -
[…]
Method additionally using images at inference time
OpenScene - OpenSeg [22] 42.1 -
AFOV [25] 47.9 -
Table 5: Semantic segmentation results on nuScenes and SemanticKITTI validation sets. $\dagger$ indicates the panoptic mIoU, which, in the case of nuScenes, slightly differs ($<$ 1 pt in general) from the classical semantic segmentation mIoU otherwise reported here (see appendix).
Comparison to state-of-the-art zero-shot panoptic segmentation
nuScenes SemanticKITTI
[…]
SAL [30] 38.4 47.8 77.2 47.5 29.2 33.9 24.8 32.3 66.8 17.4 30.2 28.7
SAL-4D [32] - - - - - - 30.8 - 76.9 25.5 34.6 -
[…]
Table 6: Panoptic segmentation results on nuScenes and SemanticKITTI validation sets.
We further evaluate our method on the panoptic segmentation task, comparing it against previous annotation-free approaches in Table 6 for both nuScenes and SemanticKITTI.
[…]
5 Limitations
The spatio-temporal consolidation (TBC) assumes that voxel semantics remains constant over time. This is true for the static part of the scene, which represents most of the points. However, it may not be true for moving objects. Yet, there is no issue when an empty area is traversed by a moving object: in corresponding voxels, there are no points before the traversal, some points during the traversal, and again no points afterwards. Therefore, voxel-based voting still makes sense.
An actual issue occurs only when two different objects traverse the same empty space, and then only in one particular case: if the objects belong to different classes (e.g., car and bicycle) and if the object speed and frame rate are such that points from different objects are actually scanned within the same voxel. Given the distribution of moving object classes (a vast majority of cars), the low frame rate (2 Hz for nuScenes, 10 Hz for SemanticKITTI), and the small voxel size (10 cm), this happens rarely.
In other words, while dynamic environments can in theory break the voxel semantic consistency, they rarely do so in practice. The fact is, LOSC outperforms all other methods while the evaluation datasets do include moving objects. Moreover, if needed, time consistency for moving objects can be improved by cutting sequences into smaller sections.
6 Conclusion
Our work shows it is possible to transfer the image segmentation capability of an off-the-shelf open-vocabulary 2D-based VLM, used in a black box manner, in order to segment 3D lidar scans and largely outperform the SOTA. Besides, compared to the bells and whistles of existing approaches, our method is relatively simple and relies on very few and little-sensitive parameters.
Our approach actually is not limited to 2D VLMs. It could actually be also applied to any strong 2D semantic segmenter trained with full-supervision, alleviating the need for 3D annotation.
Appendix
We supplement the main paper with the performance on super-classes (Appendix A), classwise results (Appendix B), information about the quantity and quality of labeled points for training (Appendix C), details about prompts used in experiments (Appendix D), a detailed study concerning the use of augmentations (Appendix E), and qualitative results (Appendix F).
Appendix A Results with super-classes
Some methods like SAL [30] and LeAP [34] evaluate on broader super-classes. For instance, SAL merges car, motorcycle, truck, etc., into a single ‘vehicle’ class. In Table 7, we recall results with the original fine-grained classes and also report our results adopting the same merging strategy as in SAL [30], applying class merging during inference to predict super-class labels. (As said in the experiment section of the paper, no comparison is possible with LeAP, which does not provides its mapping from 19 classes into 11.)
[…]
nuScenes SemanticKITTI
[…]
SAL [30] - 33.9 - 52.6 28.7 52.8
LOSC (ours) 49.3 49.8 68.4 - 35.2 67.4
[…]
Appendix B Classwise results
Tables 8 and 9 respectively present classwise semantic segmentation results on nuScenes [39] and SemanticKITTI [35] val sets, evaluated with networks initially finetuned on labelings $L_{\text{tim}}$ or $L_{\text{aug}{\otimes}\text{tim}}$ (1st iter.), before iterated training (2nd and 3rd iter.).
nuScenes $L_{\text{tim}}$ $L_{\text{aug}{\otimes}\text{tim}}$ Full
Classes (default) 1st iter. 2nd iter. 3rd iter. 1st iter. 2nd iter. 3rd iter. supp.
barrier 2.8 3.2 3.6 4.3 4.0 5.2 79.2
bicycle 28.2 28.0 27.7 21.2 20.4 21.3 53.2
bus 72.5 72.0 71.0 80.2 82.2 80.5 92.5
car 63.7 65.2 66.7 64.4 66.6 67.5 88.1
construction vehicle 2.4 4.6 5.9 9.3 9.8 10.6 50.4
motorcycle 57.0 57.6 58.7 62.6 63.5 64.1 87.8
pedestrian 67.1 67.9 68.2 70.8 72.1 72.7 83.4
traffic cone 39.7 44.9 46.5 42.3 45.5 46.3 70.3
trailer 0.0 0.1 0.1 0.1 0.0 0.0 73.5
truck 57.9 59.5 59.7 56.5 57.7 58.2 84.1
driveable surface 89.4 89.6 89.6 90.5 90.4 90.1 96.9
other flat 0.0 0.0 0.0 0.0 0.0 0.0 73.3
sidewalk 48.6 49.3 49.6 50.0 50.7 51.0 75.5
terrain 59.5 60.2 60.4 61.3 61.7 61.5 75.4
manmade 74.4 75.8 76.3 74.6 76.5 77.6 89.9
vegetation 78.4 79.4 79.7 79.5 81.0 81.7 86.5
mean 46.4 47.3 47.7 48.0 48.9 49.3 78.7
Superclasses
object 5.3 6.0 6.2 6.6 6.4 7.6
vehicle 79.9 81.8 83.3 80.5 82.4 83.0
human 67.1 67.9 68.2 70.8 72.1 72.7
ground 90.1 90.6 90.8 90.5 90.9 91.0
nature 76.7 77.6 77.9 77.3 78.2 78.5
structure 74.4 75.8 76.3 74.6 76.5 77.6
mean 65.6 66.1 67.1 66.7 67.8 68.4
Table 8: Classwise semantic segmentation results (IoU%) of LOSC on nuScenes val set when finetuning a 3D network [38, 10] on labelings $L_{\text{tim}}$ or $L_{\text{aug}{\otimes}\text{tim}}$, and iterating training interleaved with time-based consolidation. “Full supp.” is when training with ground-truth labels [38, 10]. Best results are in blue.
These tables detail the results reported in Table 3 of the main paper. They also provide classwise results for super-classes (cf. Appendix A).
SemanticKITTI $L_{\text{tim}}$ $L_{\text{aug}{\otimes}\text{tim}}$ Full
Classes (default) 1st iter. 2nd iter. 3rd iter. 1st iter. 2nd iter. 3rd iter. supp.
car 77.1 79.9 79.6 79.1 80.8 80.8 94.7
bicycle 35.6 37.4 37.5 33.3 35.5 34.7 44.7
motorcycle 58.4 61.3 63.4 58.0 56.0 56.7 64.8
truck 10.5 10.6 10.3 17.3 16.5 16.0 82.0
other-vehicle 4.8 2.7 2.3 30.8 32.3 34.1 37.4
person 52.8 54.5 55.0 55.0 56.7 56.5 71.4
bicyclist 0.0 0.0 0.0 0.0 0.0 0.0 85.9
motorcyclist 0.0 0.0 0.0 0.0 0.0 0.0 0.0
road 75.6 76.0 75.5 75.5 75.7 75.3 95.7
parking 0.0 0.0 0.0 0.0 0.0 0.0 49.2
sidewalk 55.2 55.9 55.3 55.2 55.3 54.9 83.6
other-ground 0.0 0.0 0.0 0.0 0.0 0.0 0.2
building 69.2 73.7 76.0 69.3 73.9 76.5 89.3
fence 8.0 8.9 9.3 7.3 9.2 9.0 56.0
vegetation 80.6 83.1 84.1 80.5 82.8 83.8 88.2
trunk 0.0 0.0 0.0 0.0 0.0 0.0 69.9
terrain 66.7 67.9 67.8 67.6 67.9 66.9 74.6
pole 0.0 0.0 0.0 0.0 0.0 0.0 66.0
traffic-sign 16.2 15.4 15.5 21.4 22.9 22.9 51.5
mean 32.1 33.0 33.2 34.2 35.0 35.2 63.4
Superclasses
object 7.6 8.4 8.8 7.2 9.1 8.9
vehicle 83.9 86.5 86.6 84.0 86.6 87.0
human 65.6 66.0 65.7 63.8 63.9 63.1
ground 84.0 84.3 84.0 84.3 84.1 83.6
nature 84.4 85.9 86.3 84.3 85.6 85.7
structure 69.2 73.7 76.0 69.3 73.9 76.5
mean 65.8 67.5 67.9 65.6 67.2 67.4
Table 9: Classwise semantic segmentation results (IoU%) of LOSC on SemanticKITTI val set when finetuning a 3D network [38, 10] on labelings $L_{\text{tim}}$ or $L_{\text{aug}{\otimes}\text{tim}}$, and iterating training interleaved with time-based consolidation. “Full supp.” is when training with ground-truth labels [38, 10]. Best results (besides full supervision) for each labeling are with a blue background.
For nuScenes, while the IoUof most classes keeps on increasing as we iterate more, only a few classes see their performance stagnate or slightly drop. Therefore, the overall performance (mIoU%) is clearly increasing even after 3 iterations.
For SemanticKITTI, where the overall performance is lower and increasing at a lower rate as we iterate, compared to nuScenes, a higher number of classes have a best performance at the 1st or 2nd iteration. Still, the overall performance after the 3rd iteration remains the best one.
The results for the method we call LOSC (Table 5 of the main paper) correspond to three iterations.
Appendix C Quantity and quality of labeled points for training
Table 10 below lists the number of labeled points per class in the nuScenes training set for the different labelings reported in Table 4 of the main paper. Colors in these tables illustrate the choices made in Eq. (5) to select labels that are more likely to be trustworthy, among labels from $L_{\text{tim}}$ or $L_{\text{aug}}$.
Class $L_{\text{vlm}}$ $L_{\text{tim}}$ $L_{\text{abc}}$ $L_{\text{aug}}$ $N_{\text{aug}}/N_{\text{tim}}$ $L_{\text{aug}{\otimes}\text{tim}}$ FoVgt   1st iter. 2nd iter. 3rd iter.
barrier 2.0 2.1 0.5 0.4 0.20 2.1 7.7 4.8 4.8 4.7
bicycle 0.3 0.3 0.1 0.1 0.46 0.3 0.1 0.5 0.6 0.6
bus 3.7 3.7 2.5 2.4 0.65 2.4 3.8 8.1 7.7 7.9
car 27.4 27.5 24.4 24.0 0.87 24.0 32.0 93.8 102.2 103.7
construction vehicle 0.1 0.1 0.0 0.0 0.10 0.1 1.3 0.4 0.3 0.3
motorcycle 0.3 0.3 0.2 0.2 0.70 0.3 0.3 0.3 0.3 0.3
pedestrian 2.3 2.2 1.8 1.7 0.75 1.7 1.9 1.8 1.7 1.7
traffic cone 0.2 0.2 0.1 0.1 0.45 0.2 0.6 0.5 0.5 0.5
trailer 0.2 0.1 0.0 0.0 0.06 0.1 4.5 0.9 1.1 1.2
truck 20.4 20.4 15.7 15.5 0.76 15.5 13.3 58.7 54.8 49.6
driveable surface 156.3 149.4 142.1 134.3 0.90 134.3 260.9 298.0 299.6 298.9
other flat 0.1 0.1 0.1 0.1 0.47 0.1 6.6 0.7 0.3 0.3
sidewalk 51.7 50.0 38.4 36.5 0.73 36.5 57.6 69.5 68.2 68.1
terrain 38.6 37.7 28.1 27.0 0.72 27.0 56.7 54.1 50.1 47.5
manmade 128.6 128.5 109.9 108.7 0.85 108.7 146.5 168.9 152.2 147.0
vegetation 94.8 92.4 79.0 77.1 0.83 77.1 100.6 215.8 232.3 244.4
Table 10: Quantity of labeled points used in nuScenes train set. The number of points is given in millions. Values under 50k appear as “0.0”. Values not meeting the conditions in Eq. (4) are indicated with a yellow background. Labels from $L_{\text{tim}}$ or $L_{\text{aug}}$ that are selected in $L_{\text{aug}{\otimes}\text{tim}}$ as specified in Eq. (5) are shown with a green background. Column “FoVgt” gives the number of points in the field of view of the cameras for each class according to the ground truth, for comparison purposes. The following columns (1st to 3rd iter.) provide results of trained 3D networks and thus concern all scanned points, not just those visible from the cameras.
Table 11 provides the classwise quality (IoU%) of labeled points in the nuScenes training set for the different labelings reported in Table 4 of the main paper. As in Table 10, color in this table illustrates the choices made in Eq. (5) to select labels from $L_{\text{tim}}$ or $L_{\text{aug}}$ that are more likely to be trustworthy. Note that points that receive no label with our methods are excluded when computing the metric. Hence, from one column to another (different labelings), the set of points on which the IoU is computed differs.
Class $L_{\text{vlm}}$ $L_{\text{tim}}$ $L_{\text{abc}}$ $L_{\text{aug}}$ $L_{\text{aug}{\otimes}\text{tim}}$ 1st iter. 2nd iter. 3rd iter. F. supp.
barrier 4.2 4.6 2.2 2.1 5.4 5.0 4.8 4.8 79.2
bicycle 21.0 22.0 26.5 28.0 23.6 17.8 15.7 15.9 53.2
bus 57.5 57.5 74.4 74.9 74.9 66.4 67.2 67.7 92.5
car 63.4 63.5 69.8 70.5 70.4 73.3 76.1 77.3 88.1
construction vehicle 5.4 5.2 1.4 1.1 7.2 11.9 12.1 12.4 50.4
motorcycle 41.8 43.0 50.5 52.2 48.5 57.2 59.1 59.6 87.8
pedestrian 44.0 47.0 52.5 56.6 56.4 70.5 71.7 72.0 83.4
traffic cone 24.3 26.0 25.8 26.8 34.6 45.5 48.0 48.8 70.3
trailer 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 73.5
truck 41.3 41.5 48.4 49.2 49.1 47.7 48.4 48.7 84.1
driveable surface 84.4 84.2 90.7 91.0 91.0 89.8 89.8 89.7 96.9
other flat 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 73.3
sidewalk 48.4 48.6 57.9 59.0 59.0 53.5 54.3 54.8 75.5
terrain 52.3 52.6 65.2 66.6 66.4 57.0 57.5 57.5 75.4
manmade 69.5 70.3 77.8 79.1 78.3 74.2 75.7 76.5 89.9
vegetation 71.4 72.1 79.8 81.0 81.0 79.3 80.6 81.2 86.5
mean 39.3 39.9 45.1 46.1 46.6 46.8 47.6 47.9 78.7
Table 11: Quality (IoU%) of labeled points used in nuScenes train set. Labels from $L_{\text{tim}}$ or $L_{\text{aug}}$ selected in $L_{\text{aug}{\otimes}\text{tim}}$ are shown with a green background. Columns “1st iter.” to “3rd iter.” provide results of trained 3D networks and thus concern all scanned points, not just those visible from the cameras as in the previous columns. “F. supp.” is when training with ground-truth labels [38, 10].
Appendix D Text prompts
In this section, we provide detailed class-specific prompts for both our minimal and rich text prompt settings. Table 12 lists the minimal prompts employed in our final models for nuScenes and SemanticKITTI. Additionally, Table 13 presents the class-specific rich prompts used for nuScenes in the benchmarking of 2D open-vocabulary VLMs, as reported in Table 1 of the main paper.
Table 12: Our classwise minimal text prompts used to obtain segmentation maps from OpenSeed [19].
Class Text prompts
nuScenes
pedestrian person, pedestrian
bicycle bicycle
bus bus
car car
construction vehicle bulldozer, excavator, concrete mixer , crane, dump truck
motorcycle motorcycle
trailer trailer, sem trailer, cargo container, shipping container, freight container
truck truck
barrier barrier, barricade
traffic cone traffic cone
driveable surface road
other flat curb, traffic island, traffic median
sidewalk sidewalk
terrain terrain, grass, grassland, lawn, meadow, turf, sod
manmade building, wall, pole, awning
vegetation tree, trunk, tree trunk, bush, shrub, plant, flower, woods
SemanticKITTI
car car
bicycle bicycle
motorcycle motorcycle
truck truck
other vehicle trailer, semi trailer, cargo container, shipping container, freight container, caravan, bus, bulldozer, excavator, concrete mixer, crane, dump truck, train, tram
person person, pedestrian
bicyclist bicyclist, cyclist
motorcyclist motorcyclist
road road
parking parking, parking lot
sidewalk sidewalk, curb, bike path, walkway, pavement,footpath,footway,boardwalk, driveway
other ground water, river, lake, watercourse, waterway, canal, ditch, rail track, traffic island, traffic median, median strip, roadway median, central reservation
building building, house, garage, wall, railing, stairs, awning, roof, bridge
fence fence, barrier, barricade
vegetation tree, bush, shrub, plant, flower
trunk tree trunk, trunk, woods
terrain terrain, grass, soil, grassland, hill, sand, gravel, lawn, meadow, garden, earth, peeble, rock
pole pole
traffic sign traffic-sign
Table 13: Our rich text prompts used to obtain segmentation maps from OpenSeed [19] on nuScenes.
Class Text prompts
nuScenes
pedestrian pedestrian, person, human, man, woman, adult, child, stroller, wheelchair
bicycle bicycle
bus bus, autobus, motorbus, omnibus, double-decker, jitney, minibus, motor coach, school bus, tour bus, sightseeing bus, shuttle bus, shuttle, bendy bus
car car, automobile, van, SUV, sedan, hatchback, wagon, minivan, convertible, jeep,ambulance
construction vehicle bulldozer, excavator, concrete mixer, crane, dump truck, dozer, digger, dumper, tipper truck, tipper lorry, front loader, loader, backhoe, trencher, compactor, forklift
motorcycle motorcycle, motorbike, choper, motor scooter, scooter, vespa
trailer trailer, semi trailer, cargo container, shipping container, freight container
truck truck, pickup, pickup truck, lorry, autotruck, semi, semi-tractor, semi-truck, motortruck
barrier barrier, jersey barrier, barricade
traffic cone traffic cone, safety pylon, road cone, highway cone, safety cone, caution cone, channelizing device, construction cone, cone, road delineator
driveable surface road, parking, parking lot, highway, expressway, freeway, thoroughfare, route, thruway, turnpike, roadway, macadam, lane
other flat water, river, lake, watercourse, waterway, canal, ditch, railway track, traffic island, traffic median, median strip, roadway median, central reservation
sidewalk sidewalk, curb, bike path, walkway, pavement, footpath, footway, boardwalk
terrain terrain, grass, grassland, hill, soil, sand, gravel, lawn, meadow, garden, earth, peeble, rock
manmade building, skyscraper, house, wall, stairs, awning, roof, pole, streetlight, traffic light, traffic sign, bench, fire hydrant, hydrant, guard rail, guardrails, bollard, fence, railing, drainage, flag, banner, street sign, circuit box, traffic signal, stoplight, lamppost, bus shelter, bus stop, bin, mailbox, newspaper box, newsbox
vegetation tree, bush, plant, shrub, potted plant, hedge, branch, tree trunk, flower, woods, forest
Appendix E Augmentations
The following Table 15 presents detailed classwise results of zero-shot OpenSeeD [19] applied to the nuScenes dataset, before and after using our set of 10 distinct data augmentation methods. We observe that the results remain largely consistent with the original, which helps us eliminate obviously incorrect labels. We also present the results of several other data augmentations in Table 15. These augmentations perform significantly worse than the no-augmentation baseline and result in a substantial loss of points.
Class no horizontal hue blur color auto- sharpen chromatic emboss fancy clahe
aug. flip saturation jitter contrast aberration pca
barrier 1.6 1.9 2.1 1.7 1.7 2.6 1.7 1.6 1.7 1.6 1.4
bicycle 24.3 24.2 23.9 23.5 24.3 24.0 24.2 23.8 23.7 24.3 24.0
bus 56.7 57.2 54.6 57.4 56.2 49.7 53.0 56.0 56.6 56.8 55.6
car 47.9 48.0 48.0 48.4 48.1 46.3 47.8 47.7 47.2 47.9 48.3
construction vehicle 5.2 5.1 4.0 3.7 4.2 3.1 3.7 3.3 3.8 5.2 4.2
motorcycle 31.2 31.2 30.7 31.2 30.8 29.9 30.8 30.6 31.1 31.2 29.9
pedestrian 40.4 40.5 40.3 40.6 40.6 39.5 40.2 40.3 39.7 40.4 39.0
traffic cone 20.5 20.4 19.2 21.3 20.3 22.8 18.3 19.8 19.0 20.5 16.6
trailer 0.1 0.1 0.2 0.1 0.1 0.2 0.2 0.1 0.1 0.2 0.1
truck 45.7 45.5 45.0 46.1 45.8 44.3 45.2 45.1 45.1 45.7 46.0
driveable surface 50.6 50.6 50.3 50.2 50.4 50.2 50.2 50.1 50.3 50.6 50.4
other flat 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
sidewalk 34.4 34.4 33.9 33.6 34.1 33.8 34.2 33.0 33.8 34.4 33.9
terrain 43.9 43.8 42.7 43.4 43.7 44.0 43.6 43.3 43.2 43.8 43.6
manmade 66.7 66.7 65.9 66.9 66.4 65.2 65.7 66.4 65.7 66.7 66.6
vegetation 66.7 66.6 65.7 67.4 66.2 65.6 64.4 66.8 64.9 66.7 67.7
mean IoU 33.4 33.5 32.9 33.4 33.3 32.5 32.6 32.9 32.8 33.4 32.9
mean Acc 43.4 43.4 42.8 43.3 43.1 42.8 42.4 42.7 42.4 43.4 42.9
Table 14: Classwise results on the nuScenes validation set, before and after applying our 10 selected data augmentations. These augmentations yield results comparable to the no-augmentation baseline and help eliminate unreliable predictions.
Class no salt-and gauss random random random glass shot superpixels vertical
aug. pepper noise sun-flare snow rain blur noise flip
barrier 1.6 1.6 1.0 1.3 1.3 1.1 0.5 0.9 0.9 0.0
bicycle 24.3 21.6 17.8 24.0 20.8 23.5 14.3 14.4 14.1 14.8
bus 56.7 46.9 50.0 45.5 50.5 42.9 35.2 37.3 36.8 22.8
car 47.9 48.2 48.2 45.3 48.0 46.2 47.5 45.5 39.6 36.5
construction vehicle 5.2 1.2 1.0 2.2 0.9 2.4 0.0 1.0 0.7 0.0
motorcycle 31.2 27.9 27.2 29.2 27.6 26.6 18.5 23.1 21.0 5.4
pedestrian 40.4 37.4 36.3 39.7 39.7 40.1 37.7 31.2 33.4 31.6
traffic cone 20.5 17.8 19.0 18.2 14.7 20.5 17.6 17.2 14.0 0.0
trailer 0.1 0.0 0.0 0.0 0.1 0.0 0.0 0.0 0.0 0.7
truck 45.7 43.0 45.5 42.5 44.0 39.9 37.8 35.4 35.4 22.3
driveable surface 50.6 48.4 48.7 48.9 48.7 48.2 45.1 46.6 46.6 17.1
other flat 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
sidewalk 34.4 29.8 28.9 31.0 28.8 27.4 16.5 25.2 26.4 7.5
terrain 43.9 41.0 41.6 38.9 37.6 42.6 40.0 37.3 40.5 36.6
manmade 66.7 63.2 61.3 62.7 65.3 60.1 59.9 56.3 61.4 40.3
vegetation 66.7 60.8 60.8 59.6 64.7 60.2 62.8 55.2 59.0 60.0
mean IoU 33.4 30.5 30.4 30.5 30.7 30.1 27.0 26.6 26.8 18.4
mean Acc 43.4 39.3 39.8 40.5 40.1 40.0 35.0 35.4 34.6 27.1
Table 15: Classwise results on the nuScenes validation set, comparing performance before and after applying several additional data augmentations. These augmentations significantly degraded performance compared to the baseline without any augmentation. Due to their negative impact and deviation from the original scores, we excluded them from our pipeline to ensure sufficient points in the scenes.
Appendix F Qualitative examples
Last, we present in Figure 3 qualitative results on the nuScenes and SemanticKITTI validation sets. A typical error with LOSC on both datasets is a confusion between different types of flat surfaces such as road/driveable surface, sidewalk and terrain. In SemanticKITTI, trunks are also systematically included in vegetation rather than considered as a separate class. Theses observations are consistent with the quantitative results provided in Tables 8 and 9.
Ground truth LOSC
nuScenes
Refer to caption Refer to caption
nuScenes
Refer to caption Refer to caption
SemanticKITTI
Refer to caption Refer to caption
SemanticKITTI
Refer to caption Refer to caption
Figure 3: Qualitative results of semantic segmentation from the validation sets of nuScenes and SemanticKITTI. The color code used to represent each class is provided in Fig. 4.
Color code used for nuScenes data
Color code used for SemanticKITTI data
Figure 4: Color code used to represent each class on nuScenes (top) and SemanticKIITI (bottom) in Fig. 3.